OpenStudy (anonymous):

Are the graphs of 2y = -3x + 6 and y = 2/3 x + 2, parallel, Perpendicular, or neither? How do you know?

OpenStudy (anonymous):

What do you know about perpendicular lines? What do you know about parallel lines?

OpenStudy (anonymous):

perpendicular colide parallel dont

OpenStudy (anonymous):

Correct, perpendicular lines intersect and parallel lines do not. What else do you know? Hint: slopes

OpenStudy (anonymous):

nothing rlly

jimthompson5910 (jim_thompson5910):

If two lines are perpendicular, then their slopes multiply to ____

OpenStudy (anonymous):

Are the slopes of parallel lines the same?

OpenStudy (anonymous):

no?

jimthompson5910 (jim_thompson5910):

that's incorrect, if any two lines are parallel, then they will have the same slope

OpenStudy (anonymous):

ok then what?

jimthompson5910 (jim_thompson5910):

solve 2y = -3x + 6 for y and tell me what you get

OpenStudy (anonymous):

y= -3/2x +

OpenStudy (anonymous):

3

jimthompson5910 (jim_thompson5910):

the slope of y = -3/2x + 3 is -3/2

jimthompson5910 (jim_thompson5910):

the slope of y = 2/3 x + 2 is 2/3

jimthompson5910 (jim_thompson5910):

multiply these two slopes to get (-3/2)*(2/3) = (-3*2)/(2*3) = -6/6 = -1 Since you get a product of -1, this tells us that the two lines are perpendicular

OpenStudy (anonymous):

kk